Logiweb codex of factorial

Up Help factorial, ∗!,

factorial

[factorial →^{pyk} “factorial”]

∗!

[x! →^{val} [ x ≈ 0 ] \{ \begin{matrix}1 \\ x · [ [ x − 1 ] ! ] \end{matrix} ]

[x! →^{tex} “#1.!”]

[x! →^{pyk} “∗ factorial”] The pyk compiler, version 0.grue.20050502 by Klaus Grue GRD-2005-06-02.UTC:14:20:27.677282 = MJD-53523.TAI:14:20:59.677282 = LGT-4624438859677282e-6